A bag contains 3 blue balls,5 red balls, and 4 yellow balls. One ball is taken at random,its colour noted,and then put back into the bag.A second ball is taken at random,and again its colour is noted.
What is the probability that the both balls are the same colour?

Answer by jrfrunner(365)

total number of balls= 3+5+4=12
--
P(selecting two balls with replacement of the same color)
= P(RR)+P(BB)+P(YY)=(3/12)*(3/12)+(5/12)*(5/12)+(4/12)*(4/12)
=(1/4)^2 + (5/12)^2 + (1/3)^2=1/16+25/144+1/9=0.347